OpenStudy (anonymous):

The color of some clothing fades over time when washed. Suppose a pair of jeans fades by 5% with each washing. Question: How many washings would it take so that only 25% of the original color remains in the jeans?

OpenStudy (anonymous):

\[25 = 100(0.95)^{n-1}\]

OpenStudy (anonymous):

\[.25=(.95)^x\] solve for x

OpenStudy (anonymous):

I don't get it, I need to solve for n

OpenStudy (amistre64):

\[f_n=f_{n-1}-f_{n-1}(.05)\] \[f_n=f_{n-1}(1-.05)\] \[f_n=f_{n-1}(.95)\] \[f_n=f_0(.95)^n\]

OpenStudy (anonymous):

Still doesn't make sense to me.

OpenStudy (anonymous):

Can you show steps of manipulation?

OpenStudy (amistre64):

those are the steps .... where do you get lost in them?

OpenStudy (amistre64):

since we do not have your material to guide us; we just got to take a blind stab at it

OpenStudy (anonymous):

But i did give you info, it is a geometric sequence question

OpenStudy (anonymous):

\[T_{n}=T _{1}(r)^{n-1}\]

OpenStudy (amistre64):

thats what we got; but my books always started at n=0, not n=1

OpenStudy (anonymous):

\[25 = 100(0.95)^{n-1}\]

OpenStudy (amistre64):

divide off the 100

OpenStudy (anonymous):

\[0.025 = (0.0095)^{n-1}\]

OpenStudy (amistre64):

dont add the superflous zeros

OpenStudy (amistre64):

25/100 = .95^(n-1) .25 = .95^(n-1) which is what satellite started out with

OpenStudy (amistre64):

ln both side ln.25 = ln(.95^(n-1)) ln.25 = (n-1) ln(.95) ln.25/ln(.95) = n-1